OpenStudy (anonymous):

?? Simplify completely quantity 6 x squared minus 54 plus 84 over quantity 8 x squared minus 40 x plus 48 divided by quantity x squared plus x minus 56 over quantity 2 x squared plus 12 x minus 32 2 open parentheses x minus 3 close parentheses over 3 open parentheses x minus 2 close parentheses 2 open parentheses x minus 2 close parentheses over 3 open parentheses x minus 3 close parentheses 3 open parentheses x minus 3 close parentheses over 2 open parentheses x minus 2 close parentheses 3 open parentheses x minus 2 close parentheses over 2 open parentheses x minus 3 close par

OpenStudy (anonymous):

please try to write the question in mathematical form

OpenStudy (anonymous):

\[\frac{ 6x^2-54+84 }{ 8x^2-40x+48 } \div \frac{ x^2+x-56 }{ 2x^2\div12x-32 }\]

OpenStudy (anonymous):

That's the equation to the question.

OpenStudy (anonymous):

okay, at first we factorize all the quadratic equations

OpenStudy (anonymous):

meaning?

OpenStudy (anonymous):

simplify the quadratic equations, like 6x^2−54x+84 = 6(x-7)(x-2)

OpenStudy (anonymous):

8x^2−40x+48 = 8(x-3)(x-2)

OpenStudy (anonymous):

similarly , x^2+x−56 = (x-7)(x+8)

OpenStudy (anonymous):

and, 2x2+12x−32 = 2(x-2)(x+8)

OpenStudy (anonymous):

I hope, you can do it from here